# The $\alpha$–regression for compositional data: a unified framework for standard, spatially-lagged, spatial autoregressive and geographically-weighted regression models

### v2

@@ B.1 The $\alpha$–transformation
B.2 First Derivatives (Gradient)
B.2.1 Main Gradient Formula
@@
B.2.2 Expanded Gradient Formula
@@
B.2.3 Jacobian of Power Transformation
@@
B.2.4 Jacobian of Multinomial Logit
@@
B.2.5 Vectorized Gradient Formula
@@
B.3 Hessian matrix for the $\alpha$–regression
@@
B.3.1 First-Order Term (Gauss-Newton Part)
@@
B.3.2 Second-Order Term (Exact Correction)
@@
B.3.3 Chain Rule for Second Derivative
@@
B.3.4 Second Derivative of Power Transformation
@@
B.3.5 General Formula for Hessian of Power Transformation
@@
B.3.6 Second Derivative of Multinomial Logit
@@
B.3.7 Assembling the Second-Order Term
@@
B.3.8 Complete Hessian (Exact)
@@
B.4 Gradient vecctor and Hessian matrix for the $\alpha$–SAR model
@@
Appendix C Model Specification Review
@@
C.1 First Derivatives (Gradient)
C.1.1 Gradient with respect to $\mathbf{\beta}_{k}$
@@
C.1.2 Gradient with respect to $\rho$
@@
C.2 Second Derivatives: $\mathbf{H}_{\mathbf{\beta}\mathbf{\beta}}$
@@
C.2.1 Gauss-Newton Approximation (First Term)
@@
C.2.2 Exact Correction (Second Term)
@@
C.2.3 Second Derivative of Power Transformation
@@
C.2.4 Second Derivative of Multinomial Logit
@@
C.2.5 Mixed Derivatives: $\mathbf{H}_{\mathbf{\beta}\rho}$
@@
C.2.6 Gauss-Newton Part
@@ C.2.7 Exact Correction
C.2.8 Computing $\frac{\partial^{2}\mu_{ip}}{\partial\mathbf{\beta}_{k}\partial\rho}$
@@ C.2.9 Second Derivative: $H_{\rho\rho}$
C.2.10 Gauss-Newton Part
@@ C.2.11 Exact Correction
C.2.12 Computing $\frac{\partial^{2}\mu_{ip}}{\partial\rho^{2}}$
@@
C.2.13 Complete Hessian Matrix
@@ C.2.14 Simplified Structure
Appendix D Standard error of the SMEs of the $\alpha$–SLX model
@@
Appendix E Standard error of the SMEs of the $\alpha$–SAR model

### v3

@@ D.1 The $\alpha$–transformation
D.2 First Derivatives (Gradient)
D.2.1 Main Gradient Formula
@@
D.2.2 Expanded Gradient Formula
@@
D.2.3 Jacobian of Power Transformation
@@
D.2.4 Jacobian of Multinomial Logit
@@
D.2.5 Vectorized Gradient Formula
@@
D.3 Hessian matrix for the $\alpha$–regression
@@
D.3.1 First-Order Term (Gauss-Newton Part)
@@
D.3.2 Second-Order Term (Exact Correction)
@@
D.3.3 Chain Rule for Second Derivative
@@
D.3.4 Second Derivative of Power Transformation
@@
D.3.5 General Formula for Hessian of Power Transformation
@@
D.3.6 Second Derivative of Multinomial Logit
@@
D.3.7 Assembling the Second-Order Term
@@
D.3.8 Complete Hessian (Exact)
@@
Appendix E Gradient vecctor and Hessian matrix for the $\alpha$–SAR model
@@
E.1 Model Specification Review
@@
E.2 First Derivatives (Gradient)
E.2.1 Gradient with respect to $\mathbf{\beta}_{k}$
@@
E.2.2 Gradient with respect to $\rho$
@@
E.3 Second Derivatives: $\mathbf{H}_{\mathbf{\beta}\mathbf{\beta}}$
@@
E.3.1 Gauss-Newton Approximation (First Term)
@@
E.3.2 Exact Correction (Second Term)
@@
E.3.3 Second Derivative of Power Transformation
@@
E.3.4 Second Derivative of Multinomial Logit
@@
E.3.5 Mixed Derivatives: $\mathbf{H}_{\mathbf{\beta}\rho}$
@@
E.3.6 Gauss-Newton Part
@@ E.3.7 Exact Correction
E.3.8 Computing $\frac{\partial^{2}\mu_{ip}}{\partial\mathbf{\beta}_{k}\partial\rho}$
@@ E.3.9 Second Derivative: $H_{\rho\rho}$
E.3.10 Gauss-Newton Part
@@ E.3.11 Exact Correction
E.3.12 Computing $\frac{\partial^{2}\mu_{ip}}{\partial\rho^{2}}$
@@
E.3.13 Complete Hessian Matrix
@@ E.3.14 Simplified Structure
Appendix F Standard error of the SMEs of the $\alpha$–SLX model
@@
Appendix G Standard error of the SMEs of the $\alpha$–SAR model